OpenStudy (anonymous):

Prove: tan2 θ•cos2 θ + cos2 θ = 1. I can not figure out how to answer this question! please help!

OpenStudy (anonymous):

\[\tan^2 \Theta = \frac{ \sin^2 \Theta }{ \cos^2 \Theta }\] \[\tan^2 \Theta * \cos^2 \Theta = \sin^2 \Theta\] \[\sin^2 \Theta + \cos^2 \Theta = 1\] 1 = 1